OpenStudy (anonymous):

Line OP has an equation of a line y = 6x - 5, and line QR has an equation of a line y = 6x + 3. These two equations represent Select one: a. the same line b. lines that are neither parallel nor perpendicular c. parallel lines because the slopes of the lines are equal d. perpendicular lines because the slopes are opposite reciprocals

OpenStudy (abmon98):

c) slopes are equal

OpenStudy (abmon98):

because they share the same gradient which is 6

OpenStudy (anonymous):

Look at this . this is the form for graph equations : y=mx+d m is slope y=mx+d y=6x + 3 m=6 d=3 ________________________________ y =6x-5 m=6 d=-5 Now we can say we have : equal slopes and parallel lines ;)